# Advances on the finite element discretization of fluid-structure interaction problems

### v1

@@ 1.3. How to deal with the coupling term
As we previously mentioned, the main idea behind the fictitious domain consists in extending the domain $\Omega_{1}$ to the region occupied by $\Omega_{2}$. The domain $\Omega_{2}$ is then superimposed to the background $\Omega$. Such superimposition is mathematically enforced by means of the Lagrange multiplier $\lambda$, which is applied through the coupling bilinear form $c(\cdot,\cdot)$. We notice that, at continuous level, the bilinear form $c(\cdot,\cdot)$ has the same representation when applied to functions in $H^{1}(\Omega_{2})$ and to the restriction to $\Omega_{2}$ of elements in $H^{1}_{0}(\Omega)$. On the other hand, at discrete level, Problem 3 in matrix form reads
@@
Since the mesh $\mathcal{T}_{h}^{2}$ of $\Omega_{2}$ is used to discretize both $u_{2}$ and $\lambda$, the matrix $\mathsf{C}_{2}$ is assembled with the standard procedure of finite elements. Let us focus on the interface matrix $\mathsf{C}_{1}$. The assembly procedure for this finite element matrix is not trivial since integration over non-matching grids is required (see [18]). Indeed, the definition of $\mathsf{C}_{1}$ associated with the scalar product either in $L^{2}(\Omega_{2})$ or $H^{1}(\Omega_{2})$ requires the integration over $\Omega_{2}$ of $\zeta_{k}\in\Lambda_{h}$, which is defined on $\Omega_{2}$, multiplied by $\varphi_{i}$, which is a function defined on the background mesh $\mathcal{T}_{h}$ and then restricted to the immersed domain. An example of configuration is shown in Figure 2a: it is clear that the immersed sandy colored element is not matching with the support of the finite element basis functions defined on the background mesh.
@@
[…]
@@
[…]
We are going to discuss how to compute the local integrals reported in the above equations. We will focus on two possible techniques. Indeed, the exact computation of such quantities requires the design of a composite quadrature rule taking into account the intersection between $K_{2}\in\mathcal{T}_{h}^{2}$ and the background mesh $\mathcal{T}_{h}$. We could also use a quadrature rule directly on each $K_{2}\in\mathcal{T}_{h}^{2}$, which leads to a quadrature error. In the remainder of this section we restrict the discussion to the two-dimensional case in order to keep the presentation simpler.
@@ 1.3.1. Exact procedure
as shown in Figure 2b. More precisely, each $P_{j}$ is contained in a single element $K\in\mathcal{T}_{h}$, so that the background basis functions are polynomials in $P_{j}$. We have at least two options for integrating in each $P_{j}$. If $P_{j}$ is not a triangle or a quadrilateral, we construct an auxiliary triangulation $\{T_{i}\}_{i=1,\dots,I(j)}$ by simply connecting the barycenter with the vertices. This gives
[…]
@@ 1.3.2. Inexact procedure
The evaluation of the intersection between the involved meshes may be expensive, unless special computational techniques are employed for an efficient implementation.
@@
[…]
@@
Remark 4.
In [19], we studied the behavior of the quadrature error in the case of FSI problems discretized by piecewise linear elements on triangular meshes, however the same arguments are still valid in the case of simpler elliptic interface problems. We proved that if $c(\cdot,\cdot)$ is the scalar product in $L^{2}(\Omega_{2})$, then the quadrature error is comparable to the other discretization errors. On the other hand, if the $H^{1}(\Omega_{2})-$scalar product is considered, then the quadrature error is large unless the ratio $h_{2}/h$ is decaying fast enough. The interested reader may refer to [19] for a complete discussion about these aspects.
Remark 5.
The same computational techniques can be considered when solving three dimensional problems. We have not proved convergence estimates for the quadrature error in the case of 3D problems, but we expect the same behavior described in the previous remark.
Remark 6.
@@ 1.4. Solvers
The linear system (8) arising from Problem 3 is usually large and its direct solution is not feasible. The use of an iterative solver is thus recommended and GMRES represents a good option, although the design of effective preconditioners is not straightforward. We presented and discussed a first parallel solver for the fictitious domain formulation of fluid-structure interaction problems in [21, 22], where we performed an in depth numerical investigation to assess the robustness of the solver with respect to mesh refinement and strong scalability. In these works, we proposed a diagonal and a triangular preconditioner showing that the triangular one is better performing.
@@ 1.5. A posteriori error analysis
As discussed in Remark 1, the interface problem under consideration is characterized by solutions with reduced regularity. While quasi-optimal convergence rates can be achieved with uniform mesh refinement, mesh adaptivity remains essential for recovering optimal rates by concentrating computational effort near singularities and interfaces.
@@
We introduce four residuals: $r_{K}$, $r_{K_{2}}$ for elements, and $r_{e}$, $r_{e_{2}}$ for edges. In addition, we introduce suitable extensions $\tilde{\lambda}$ and $\tilde{\lambda}_{h}$ of the continuous and discrete Lagrange multipliers, see [4, Def. 3.1, Def. 3.2]. More precisely, we extend by zero $\lambda_{h}$ to $\Omega$. Hence, we have $\tilde{\lambda}_{h}:\Omega\rightarrow\mathbb{R}$ with $\tilde{\lambda}_{h|\Omega_{2}}=\lambda_{h}$ and $\tilde{\lambda}_{h|\Omega\setminus\Omega_{2}}=0$. The definition of the extension of $\lambda$ is given by $\tilde{\lambda}\in H^{-1}(\Omega)$ such that
@@
Here $\langle\cdot,\cdot\rangle_{W}$ denotes the duality pairing between $W$ and its dual space $W^{\prime}$. We observe that the first equation in Problem 2 reads
@@
Similarly, for any $K_{2}\in\mathcal{T}_{h}^{2}$, we have
@@
(16) $r_{e}(u_{h})=\begin{cases}-\left\llbracket\nu\nabla u_{h}\cdot\mathbf{n}_{1}\right\rrbracket_{e},&\text{if }e\text{ is an interior edge},\\ 0,&\text{if }e\text{ is on }\partial\Omega.\end{cases}$
@@
(17) $r_{e_{2}}(u_{2,h})=\begin{cases}-\left\llbracket(\nu_{2}-\nu)\nabla u_{2,h}\cdot\mathbf{n}_{2}\right\rrbracket_{e_{2}},&\text{if }e_{2}\text{ lies in theinterior of }\mathcal{T}_{h}^{2},\\ -(\nu_{2}-\nu)\nabla u_{2,h}\cdot\mathbf{n}_{2},&\text{if }e_{2}\text{ is on }\Gamma.\end{cases}$
@@
The error indicators are constructed by combining weighted norms of the residuals with the oscillation terms, yielding a comprehensive measure of the local error distribution. Then, for each element $K_{2}\in\mathcal{T}_{h}$ and $K_{2}\in\mathcal{T}^{2}_{h}$, the local error indicators are given by:
@@
As usual, the global error estimator is the aggregation of all local contributions:
@@
Proposition 3.
Let $(u,u_{2},\lambda)$ be the exact solution of the continuous problem and $(u_{h},u_{2,h},\lambda_{h})$ be the solution of the discrete one. Then, there exists a constant ${C}>0$ that depends only on the domains $\Omega$ and $\Omega_{2}$, the coefficients $\nu$ and $\nu_{2}$, and the shape regularity of $\mathcal{T}_{h}$ and $\mathcal{T}^{2}_{h}$, such that
@@
The proof, which involves the approximation of functions in $H^{1}$ by means of a Clément-type interpolation, exploits the definition of the extension of $\lambda$ given in (13), and that by zero of $\lambda_{h}$, see [4, Prop. 3.1]. This result also holds to the case of non-constant coefficients $\nu$ and $\nu_{2}$ as described in [4, Sect. 3.2].
@@
The proof of this proposition relies on the arguments in [73], which make use of bubble functions in order to localize the indicators on elements or edges. We refer to [4, Prop. 3.2] for the details of the proof in the case of constant coefficients $\nu$, $\nu_{2}$ and [4, Sect. 3.2] for the general case.
@@ 1.7.1. Preconditioning strategies
The linear system is solved using GMRES with a tolerance of $10^{-12}$. Moreover, we employ the preconditioners defined in (12). Their action requires the inversion of the two diagonal blocks. This can be done by using a direct solver, denoted by “$\mathsf{d}$”, or by a multigrid one, “$\mathsf{m}$”. We consider three configurations for inverting the blocks of the preconditioner matrix: $\mathsf{dd}$, $\mathsf{md}$, and $\mathsf{mm}$. Here, $\mathsf{md}$ means that the first block is inverted with multigrid, while $\mathsf{dd}$ and $\mathsf{mm}$ refer to the inversion of both blocks with either the direct or multigrid method.
Table 1 reports the iteration count $its$ and solution time $T_{sol}$ for the case $\Lambda=H^{1}(\Omega_{2})$ with the element $(\mathcal{Q}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$. The results indicate that the diagonal preconditioner combined with $\mathsf{dd}$ inversion performs poorly. However, the $\mathsf{md}$ and $\mathsf{mm}$ configurations are both effective in controlling the iteration count and solution time, with $\mathsf{md}$ showing a slightly better performance. For the lower triangular preconditioner, all three inversion strategies yield comparable performance.
@@
$dofs$ $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
$its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$
@@
Table 1. Mesh refinement test for the immersed circular shape, $H^{1}$ coupling term with mesh intersection and $(\mathcal{Q}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$ element. The simulations are run on one processor. Legend: $dofs$ = degrees of freedom; $its$ = GMRES iterations; $T_{sol}$ = CPU time to solve the linear system. All CPU times are reported in seconds.
@@
$dofs$ $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
$its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$
@@
$dofs$ $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
$its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$ $its$ $T_{sol}(s)$
@@ 2. Application to fluid-structure interaction problems
In this section we review how to apply our fictitious domain formulation to fluid-structure interaction problems where a viscous hyper-elastic solid body is immersed in a Newtonian fluid.
@@ 2.1. The continuous problem
We study fluid-structure interaction problems in a fixed domain $\Omega\subset\mathbb{R}^{d}$ with $d=2,3$. Such domain is decomposed into two time dependent regions, $\Omega^{f}_{t}$ and $\Omega^{s}_{t}$, occupied by the evolving fluid and solid, respectively, at time instant $t$. We assume that $\Omega^{s}_{t}$ has codimension zero, even if more general situations, such as immersed thin structures, can be modeled by our approach (see e.g. [17, 6, 5, 53] for more details). Moreover, for simplicity of exposition, we assume that the immersed interface $\Gamma_{t}=\overline{\Omega}_{t}^{s}\cap\overline{\Omega}_{t}^{f}$ cannot touch $\partial\Omega$, that is $\Gamma_{t}\cap\partial\Omega=\emptyset$.
We describe the fluid dynamics in an Eulerian framework and we denote by $\mathbf{x}$ the associated variable. The dynamics of incompressible Newtonian fluids is governed by the Navier–Stokes equations in $\Omega^{f}_{t}$. Given the viscosity $\nu_{f}>0$, the density $\rho_{f}$, the velocity $\mathbf{u}_{f}$ and the pressure $p_{f}$, the Cauchy stress tensor for viscous fluids reads
@@
The solid deformation is described in a Lagrangian setting. We consider a fixed domain $\mathcal{B}\in\mathbb{R}^{d}$ playing the role of reference domain for the solid configuration. At each time instant $t$, the actual configuration of the solid is obtained through the action of the deformation map ${\mathbf{X}(\cdot,t):\mathcal{B}\rightarrow\Omega^{s}_{t}}$. We have that
@@
where the Lagrangian variable is denoted by $\mathbf{s}\in\mathcal{B}$. An example of geometric configuration for the FSI system is shown in Figure 6.
@@
where the pressure $p_{s}$ plays the role of a Lagrange multiplier enforcing the incompressibility of the solid material. Concerning the elastic contribution, $\bm{\sigma}_{s}^{e}$ is expressed in terms of the first Piola–Kirchhoff elasticity tensor $\mathbb{P}$, that is,
@@
Figure 6. Geometric configuration of the FSI problem. The domain $\Omega$ is fixed in time and behaves as a container. The fluid and solid regions, $\Omega^{f}_{t}$ and $\Omega^{s}_{t}$ respectively, are time dependent. The Lagrangian description is used to represent the solid deformation: the reference domain $\mathcal{B}$ is mapped into the actual position $\Omega^{s}_{t}$ through the map $\mathbf{X}$.
@@
(25) $\rho_{f}\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\operatorname*{\bm{\nabla}}\mathbf{u}\right)=\operatorname{div}\bm{\sigma}_{f}$ in $\text{in }\Omega^{f}_{t}$
$\operatorname{div}\mathbf{u}_{f}=0$ in $\text{in }\Omega^{f}_{t}$
@@
$\operatorname{div}\mathbf{u}_{s}=0$ in $\text{in }\Omega^{s}_{t}$
$\mathbf{u}_{f}=\frac{\partial\mathbf{X}}{\partial t}$ on $\text{on }\Gamma_{t}$
$\bm{\sigma}_{f}\mathbf{n}_{f}=-\left(\bm{\sigma}_{s}^{v}+|\mathbb{F}|^{-1}\mathbb{P}\mathbb{F}^{\top}\right)\mathbf{n}_{s}$ on $\text{on }\Gamma_{t}.$
The equations are then completed by the no-slip condition $\mathbf{u}_{f}=0$ on $\partial\Omega$ and the initial conditions
$\mathbf{u}_{f}(0)=\mathbf{u}_{f,0}\quad\text{ in }\Omega^{f}_{0},\qquad\qquad\mathbf{u}_{s}(0)=\mathbf{u}_{s,0}\quad\text{ in }\Omega^{s}_{0},\qquad\qquad\mathbf{X}(0)=\mathbf{X}_{0}\quad\text{ in }\mathcal{B}.$
@@
This constraint is enforced at variational level by means of a distributed Lagrange multiplier and a coupling bilinear form $\mathbf{c}(\cdot,\cdot)$, both defined on a suitable functional space ${\bm{\Lambda}}$, as explained in the case of elliptic interface problems (see (2)). Also in this case, we consider ${\bm{\Lambda}}$ either as $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$, or $\mathbf{H}^{1}(\mathcal{B})$. In the former case, $\mathbf{c}(\cdot,\cdot)$ is the duality pairing between $\mathbf{H}^{1}(\mathcal{B})$ and $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$, whereas in the latter $\mathbf{c}(\cdot,\cdot)$ is defined as the scalar product in $\mathbf{H}^{1}(\mathcal{B})$.
The fluid-structure interaction problem in the fictitious domain framework is the following.
Problem 4.

### v2

@@ 1.3. How to deal with the coupling term
As we previously mentioned, the main idea behind the fictitious domain consists in extending the domain $\Omega_{1}$ to the region occupied by $\Omega_{2}$. The domain $\Omega_{2}$ is then superimposed to the background $\Omega$. Such superimposition is mathematically enforced by means of the Lagrange multiplier $\lambda$, which is applied through the coupling bilinear form $c(\cdot,\cdot)$. We notice that, at continuous level, the bilinear form $c(\cdot,\cdot)$ has the same representation when applied to functions in $H^{1}(\Omega_{2})$ and to the restriction of elements in $H^{1}_{0}(\Omega)$. On the other hand, at discrete level, Problem 1.7 in matrix form reads
@@
Since the mesh $\mathcal{T}_{h}^{2}$ of $\Omega_{2}$ is used to discretize both $u_{2}$ and $\lambda$, the matrix $\mathsf{C}_{2}$ is assembled with the standard procedure of finite elements. Hence, we focus on the interface matrix $\mathsf{C}_{1}$. The assembly procedure for this finite element matrix is not trivial since integration over non-matching grids is required (see [38]). Indeed, the definition of $\mathsf{C}_{1}$ associated with the scalar product either in $L^{2}(\Omega_{2})$ or $H^{1}(\Omega_{2})$ requires the integration over $\Omega_{2}$ of $\zeta_{k}\in\Lambda_{h}$, which is defined on $\Omega_{2}$, multiplied by $\varphi_{i}$, which is a function defined on the background mesh $\mathcal{T}_{h}$ and then restricted to the immersed domain. An example of configuration is sketched in Figure 2a: it is clear that the immersed beige element is not matching with the support of the finite element basis functions defined on the background mesh.
@@
(10) $c(\mu_{h},v_{h|\Omega_{2}})=\int_{\Omega_{2}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}=\sum_{K_{2}\in\mathcal{T}_{h}^{2}}\int_{K_{2}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2},$
@@
$=\sum_{K_{2}\in\mathcal{T}_{h}^{2}}\left[\int_{K_{2}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}+\int_{K_{2}}\operatorname{\nabla}\mu_{h}\cdot\operatorname{\nabla}v_{h|\Omega_{2}}\,dx_{2}\right].$
We are going to discuss how to compute the local integrals reported in the above equations. We will focus on two possible techniques. Indeed, the exact computation of such quantities requires the design of a composite quadrature rule on the intersection between $K_{2}\in\mathcal{T}_{h}^{2}$ and the background mesh $\mathcal{T}_{h}$. We could also use a quadrature rule directly on each $K_{2}\in\mathcal{T}_{h}^{2}$, which leads to a quadrature error. In the remainder of this section we restrict the discussion to the two-dimensional case in order to keep the presentation simpler.
@@ 1.3.1. Exact procedure
as depicted in Figure 2b. More precisely, each $P_{j}$ is contained in a single element $K\in\mathcal{T}_{h}$, so that the background basis functions are completely supported in $P_{j}$. We have at least two options for integrating in each $P_{j}$. If $P_{j}$ is not a triangle or a quadrilateral, we construct an auxiliary triangulation $\{T_{i}\}_{i=1,\dots,I(j)}$ by simply connecting the barycenter with the vertices. This gives
$\int_{K_{2}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}=\sum_{j=1}^{J}\int_{P_{j}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}$
$\hskip 76.82234pt=\sum_{j=1}^{J}\sum_{i=1}^{I(j)}\int_{T_{i}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}=\sum_{j=1}^{J}\sum_{i=1}^{I(j)}\mathrm{area}(T_{i})\sum_{\ell=1}^{L_{0}}\omega_{\ell}^{0}\,\mu_{h}(\mathbf{p}_{\ell}^{0})\cdot v_{h|\Omega_{2}}(\mathbf{p}_{\ell}^{0})$
@@ 1.3.2. Inexact procedure
The computation of the intersection between the involved meshes may be expensive, unless special computational techniques are employed for an efficient implementation.
@@
$\int_{K_{2}}\mu_{h}\cdot v_{h|\Omega_{2}}\,dx_{2}\approx\mathrm{area}(K_{2})\sum_{\ell=1}^{L_{0}}\omega_{\ell}^{0}\,\mu_{h}(\mathbf{p}_{\ell}^{0})\cdot v_{h|\Omega_{2}}(\mathbf{p}_{\ell}^{0})$
@@
Remark 1.9.
In [39], we studied the behavior of the quadrature error in the case of FSI problems discretized by piecewise linear elements on triangular meshes, however the same arguments are still valid in the case of simpler elliptic interface problems. We proved that if $c(\cdot,\cdot)$ is the scalar product in $L^{2}(\Omega_{2})$, then the quadrature error is comparable to the other discretization errors. Conversely, if the $H^{1}(\Omega_{2})$-scalar product is considered, then the quadrature error is large unless the ration $h_{2}/h$ is decaying fast enough. The interested reader may refer to [39] for a complete discussion about these aspects.
Remark 1.10.
We discussed the assembly techniques for the interface matrix in the case of two dimensional problems, where triangular or quadrilateral meshes are employed. The same computational techniques can be considered when solving three dimensional problems. We have not proved convergence estimates for the quadrature error in the case of 3D problems, but we expect the same behavior described in the previous remark.
Remark 1.11.
@@ 1.4. Solvers
The linear system (8) arising from Problem 1.7 is usually large and its direct solution is not feasible. The use of an iterative solver is thus recommended and GMRES represents a good option, although the design of effective preconditioners is not straightforward. We presented and discussed a first parallel solver for the fictitious domain formulation of fluid-structure interaction problems in [71, 45], where we performed an in depth numerical investigation to assess the robustness of the solver with respect to mesh refinement and strong scalability. In these works, we proposed a diagonal and a triangular preconditioner showing that the triangular one is better performing.
@@ 1.5. A posteriori error analysis
As discussed in Remark 1.2, the interface problem under consideration is characterized by solutions with reduced regularity. While quasi-optimal convergence rates can be achieved with uniform mesh refinement, mesh adaptivity remains essential for recovering optimal rates by concentrating computational effort near singularities and interfaces.
@@
1.5.1. Residuals and error indicators
We introduce four residuals: $r_{K}$, $r_{K_{2}}$ for elements, and $r_{e}$, $r_{e_{2}}$ for edges. In addition, we introduce $\tilde{\lambda}$ and $\tilde{\lambda}_{h}$ to represent suitable extensions of the continuous and discrete Lagrange multipliers, see [50, Def. 3.1, Def. 3.2]. More precisely, we extend by zero $\lambda_{h}$ to $\Omega$. Hence, we have $\tilde{\lambda}_{h}:\Omega\rightarrow\mathbb{R}$ with $\tilde{\lambda}_{h|\Omega_{2}}=\lambda_{h}$ and $\tilde{\lambda}_{h|\Omega\setminus\Omega_{2}}=0$. The definition of the extension of $\lambda$ is given by $\tilde{\lambda}\in H^{-1}(\Omega)$ such that
@@ 1.5.1. Residuals and error indicators
Here $\langle\cdot,\cdot\rangle_{W}$ denotes the duality pairing between $W$ and its dual space $W^{\prime}$. We observe that the first equation in Problem 1.3 reads
@@
Similarly, for any $K_{2}\in\mathcal{T}_{h}^{2}$, we define
@@
[…]
@@
[…]
@@
The error indicators are constructed by combining weighted norms of the residuals with the oscillation terms, yielding a comprehensive measure of the local error distribution. Then, for each element $K_{2}\in\mathcal{T}_{h}$ and $K_{2}\in\mathcal{T}^{2}_{h}$, the local error indicators are defined by:
@@
The global error estimator is then defined by the aggregation of all local contributions:
@@
Proposition 1.13.
Let $(u,u_{2},\lambda)$ be the exact solution of the continuous problem and $(u_{h},u_{2,h},\lambda_{h})$ be the solution of the discrete one. Then, there exists a constant C>0{C}>0 that depends only on the domains $\Omega$ and $\Omega_{2}$, the coefficients $\nu$ and $\nu_{2}$, and the shape regularity of $\mathcal{T}_{h}$ and $\mathcal{T}^{2}_{h}$, such that
@@
The proof, which involve the approximation of functions in $H^{1}$ by means of a Clément-type interpolation, exploits the definition of the extension of $\lambda$ given in (13), and that by zero of $\lambda_{h}$, see [50, Prop. 3.1]. This result also holds to the case of non-constant coefficients $\nu$ and $\nu_{2}$ as described in [50, Sect. 3.2].
@@
The proof of this proposition relies on the arguments in [73], which make use of bubble functions in order to localize the indicators on elements or edges. We refer to [50, Prop. 3.2] for the details of the proof in the case of constant coefficients $\nu$, $\nu_{2}$ and [50, Sect. 3.2] for the general case.
@@ 1.7.1. Preconditioning strategies
The linear system is solved using GMRES with a tolerance of $10^{-12}$. Moreover, we employ the preconditioners defined in (12). Their action requires the inversion of the two diagonal blocks. This can be done by using a direct solver, denoted by $\mathsf{d}$”, or by a multigrid one, $\mathsf{m}$”. We consider three configurations for inverting the blocks of the preconditioner matrix: $\mathsf{dd}$, $\mathsf{md}$, and $\mathsf{mm}$. Here, $\mathsf{md}$ means that the first block is inverted with multigrid, while $\mathsf{dd}$ and $\mathsf{mm}$ refer to the inversion of both blocks with either the direct or multigrid method.
Table 1 reports the iteration count itsits and solution time $T_{sol}$ for the case $\Lambda=H^{1}(\Omega_{2})$ with the element $(\mathcal{Q}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$. The results indicate that the diagonal preconditioner combined with $\mathsf{dd}$ inversion performs poorly. However, the $\mathsf{md}$ and $\mathsf{mm}$ configurations are both effective in controlling the iteration count and solution time, with $\mathsf{md}$ showing a slightly better performance. For the lower triangular preconditioner, all three inversion strategies yield comparable performance.
@@
dofsdofs $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$
@@
Table 1. Mesh refinement test for the immersed circular shape, $H^{1}$ coupling term with mesh intersection and $(\mathcal{Q}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$ element. The simulations are run on one processor. Legend: dofsdofs = degrees of freedom; itsits = GMRES iterations; $T_{sol}$ = CPU time to solve the linear system. All CPU times are reported in seconds.
@@
dofsdofs $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$
@@
dofsdofs $\mathsf{P}_{\mathsf{diag}}$ $\mathsf{P}_{\mathsf{tri}}$
@@
itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$ itsits $T_{sol}(s)$
@@ 2. Application to fluid-structure interaction problems
In this section we apply our fictitious domain formulation for modeling fluid-structure interaction problems where a viscous hyper-elastic solid body is immersed in a Newtonian fluid.
@@ 2.1. The continuous problem
We study fluid-structure interaction problems in a fixed domain $\Omega\subset\mathbb{R}^{d}$ with $d=2,3$. Such domain is decomposed into two time dependent regions, $\Omega^{f}_{t}$ and $\Omega^{s}_{t}$, occupied by the evolving fluid and solid, respectively, at time instant $t$. We assume that $\Omega^{f}_{t}$ and $\Omega^{s}_{t}$ have codimension zero, even if more general situations, such as immersed thin structures, can be modeled by our approach (see e.g. [37, 75, 76, 77] for more details). Moreover, for simplicity of exposition, we assume that the immersed interface $\Gamma_{t}=\overline{\Omega}_{t}^{s}\cap\overline{\Omega}_{t}^{f}$ cannot touch $\partial\Omega$, that is $\Gamma_{t}\cap\partial\Omega=\emptyset$.
We describe the fluid dynamics in Eulerian framework and we denote by $\mathbf{x}$ the associated variable. The dynamics of incompressible Newtonian fluids is governed by the Navier–Stokes equations in $\Omega^{f}_{t}$. Given the viscosity $\nu_{f}>0$, the density $\rho_{f}$, the velocity $\mathbf{u}_{f}$ and the pressure $p_{f}$, the Cauchy stress tensor for viscous fluids reads
@@
The solid deformation is described in Lagrangian setting. We consider a fixed domain $\mathcal{B}\in\mathbb{R}^{d}$ playing the role of reference domain for the solid configuration. At each time instant $t$, the actual configuration of the solid is then obtained through the action of the deformation map $\mathbf{X}(\cdot,t):\mathcal{B}\rightarrow\Omega^{s}_{t}$}. We have that
@@
where the Lagrangian variable is denoted by $\mathbf{s}\in\mathcal{B}$. An example of geometric configuration for the FSI system is sketched in Figure 6.
@@
where the pressure $p_{s}$ plays the role of Lagrange multiplier to enforce the incompressibility of the solid material. Concerning the elastic contribution, $\bm{\sigma}_{s}^{e}$ is expressed in terms of the first Piola–Kirchhoff elasticity tensor $\mathbb{P}$, which gives
@@
Figure 6. Geometric configuration of the FSI problem. The domain $\Omega$ is fixed in time and behaves as a container. The fluid and solid regions, $\Omega^{f}_{t}$ and $\Omega^{s}_{t}$ respectively, are time dependent. The Lagrangian description is employed to represent the solid deformation: the reference domain $\mathcal{B}$ is mapped into the actual position $\Omega^{s}_{t}$ through the map $\mathbf{X}$.
@@
$\rho_{f}\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\operatorname*{\bm{\nabla}}\mathbf{u}\right)=\operatorname{div}\bm{\sigma}_{f}$ $\text{in }\Omega^{f}_{t}$
$\operatorname{div}\mathbf{u}_{f}=0$ $\text{in }\Omega^{f}_{t}$
@@
$\operatorname{div}\mathbf{u}_{s}=0$ $\text{in }\Omega^{s}_{t}$
$\mathbf{u}_{f}=\frac{\partial\mathbf{X}}{\partial t}$ $\text{on }\Gamma_{t}$
$\bm{\sigma}_{f}\mathbf{n}_{f}=-\left(\bm{\sigma}_{s}^{v}+|\mathbb{F}|^{-1}\mathbb{P}\mathbb{F}^{\top}\right)\mathbf{n}_{s}$ $\text{on }\Gamma_{t}.$
The model is then completed by adding the no-slip condition $\mathbf{u}_{f}=0$ on $\partial\Omega$ and the initial conditions
[…]
@@
This constraint is enforced at variational level by means of a distributed Lagrange multiplier and a coupling bilinear form $\mathbf{c}(\cdot,\cdot)$, both defined on a suitable functional space $\bm{\Lambda}$}, as explained in the case of elliptic interface problems (see (2)). Also in this case, $\bm{\Lambda}$} is defined either as $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$, namely the dual space of $\mathbf{H}^{1}(\mathcal{B})$, or $\mathbf{H}^{1}(\mathcal{B})$ itself. In the former case, $\mathbf{c}(\cdot,\cdot)$ is the duality pairing between $\mathbf{H}^{1}(\mathcal{B})$ and $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$, whereas in the latter $\mathbf{c}(\cdot,\cdot)$ is defined as the scalar product in $\mathbf{H}^{1}(\mathcal{B})$.
The fluid-structure interaction problem in fictitious domain framework is the following.
Problem 2.1.